# Do Vision & Language Decoders use Images and Text equally? How Self-consistent are their Explanations?

### v2

@@ 6.4 Closed Source Models
7 Ethics Statement
The publicly available datasets and models used in this paper may inherently include their biases and imperfections Meister et al. (2022); Garcia et al. (2023). Nonetheless, the method introduced in this paper facilitates model interpretation, which could assist future research in identifying and addressing potentially harmful model properties, behaviors, and biases.
Appendix A Appendix
A.1 Models
@@ A.2 Prompts
A.3 Additional Results with VLMs
@@
LV-Mistral 54 52 44 100
LV-Vicuna 54 66 54 100
T-SHAP pred. (%) BakLLaVA 87 90 86 88
@@
CoT Accuracy (%) BakLLaVA 11 5 8 77
LV-Mistral 4 10 8 64
LV-Vicuna 14 14 14 65
@@
Table 2: Performance, MM scores, and self-consistency scores (post-hoc and CoT explanation settings) of three VL models on data from VQA, GQA, GQA balanced (generative tasks), and MSCOCO (pairwise multiple-choice) on 100 samples each.
@@
LV-Mistral 98 80 82 86 66 90 85 88 88 80 98 85$\pm$9
LV-Vicuna 94 80 82 78 60 88 78 82 78 80 98 82$\pm$10
T-SHAP pred. (%) BakLLaVA 89 88 88 87 88 88 88 88 87 87 88 88$\pm$1
@@
CoT $acc_{r}$ (%) 50% random baseline BakLLaVA 65 65 59 64 54 57 60 57 49 58 62 59$\pm$5
LV-Mistral 70 60 56 50 56 62 58 60 60 50 42 57$\pm$7
LV-Vicuna 58 58 54 66 58 55 72 56 60 65 58 60$\pm$5
@@
Figure 7: Standard deviations for accuracy and T-SHAP over three runs for the existence instrument (pairwise multiple-choice setting) on the left and VQA (generative setting) on the right. Note: The T-SHAP plots on the right and left are not exactly identical, but the results are so similar between VQA and existence, that the plots look the same.
@@
A.4 Examples of Test Results on Individual Instances for VLMs
@@
For the CC-SHAP examples, we also show the MM-SHAP values for prediction and explanation, respectively.

### v4

@@ A.1 Closed Source Models
A.2 Compute Requirements
Future work might be interested in reducing the compute requirements solicited by self-consistency tests. CC-SHAP needs ~7 minutes to run an example on a NVIDIA A40 48 GB GPU, and the other tests need ~1.5 minutes. CC-SHAP requires ~25 minutes to compute self-consistency per sample on a NVIDIA A40 48 GB GPU for LLaVA-NeXT models running with FlashAttention-2 (Dao, 2023) and quantisation (Dettmers et al., 2023) (otherwise the memory requirements exceed 48 GB), while the other tests need ~7 minutes. Since they need to process image tokens next to text tokens, VLMs have longer input sequences compared to LLMs. LLaVA-NeXT variants even more so, because they process the image at five resolutions (increasing the image sequence length by 5×5$\times$5 ×). CC-SHAP runs multiple model inferences to estimate Shapley values, while the other self-consistency tests require at most two model inferences (e.g., Biasing Features Turpin et al., 2023). We think that CC-SHAP’s compute time is well invested, since it is a more effective measure: it adds interpretability and delivers better scores in settings such as VQA, where other tests deliver contradictory results ( either 0% or 100%, cf. Tab. 2).
A.3 Further Investigations on Methodology
While the normalisation step described in Section 3 is needed to compare values that otherwise would have different ranges, the question of aggregation is more open. Averaging is a fair strategy because we expect all tokens to be equally important. At present, we adopted this strategy, without being able to certify its superiority relative to other options. In our experiments it seemed to work out rather well. However, future work may consider and investigate other options.
Appendix B Appendix B
B.1 Background on Shapley Values
Shapley values originate from game theory (Shapley, 1953) to estimate fair rewards among players in a cooperative game. In machine learning, the game outcome is the model’s prediction and the players are parts of the input (features or tokens). In SHAP (Lundberg & Lee, 2017), the input features or tokens are assigned Shapley values that represent their importance player to the model’s prediction.
More formally, for a transformer input consisting of $p$ tokens $\{1,2,...j,...,p\}$, we form subsets $S\subseteq\{1,\ldots,p\}$ of tokens representing a coalition towards the model prediction $val(S)$ (e.g., the probability for the output in a classification setting). Tokens not being part of the subset are inactivated (e.g, deleted, masked). $val(\emptyset)$ is the output of the model when all tokens are inactive. The Shapley value for a token $j$ is computed with equation equation 5:
$\phi_{j}=\sum_{S\subseteq\{1,\ldots,p\}\backslash\{j\}}\frac{val(S\cup\{j\})-val(S)}{\gamma}$ (5)
Here, $\gamma=\frac{(p-1)!}{|S|!(p-|S|-1\mid)!}$ is the normalising factor that accounts for all possible combinations of choosing subsets $S$. When masking $p$ tokens, the number $n$ of possible coalitions grows exponentially ($n=2^{p}$), therefore it is common practice to approximate Shapley values with Monte Carlo, by randomly sub-sampling only $n=2p+1$ coalitions.
The Shapley value of a token measures its contribution towards the model prediction (e.g., the probability of image-sentence alignment) and can be positive (increases the model prediction) or negative (decreases it) or zero (no effect). Shapley values exhibit four defining properties of a fair payout, which are beneficial for model interpretability:
  • Efficiency: The contributions of all players and the value of a model prediction without any input tokens $val(\emptyset)$ sum up to the model outcome.
    $val(S)=val(\emptyset)+\textstyle\sum_{j}^{p}\phi_{j}$ (6)
  • Symmetry: Any two players that contribute equally are assigned the same payout.
  • Dummy: A non-contributing part is assigned zero value.
  • Additivity enables us to simply average the Shapley Values to determine the overall player contributions in a game with combined payouts (e.g., the two halves of a soccer match, or ensembling of decision trees).
B.2 Background on CC-SHAP
In this work, we compute CC-SHAP as described by (Parcalabescu & Frank, 2024):
Contribution Ratios for outputs of length one.
We start with the base case, where the LLM predicts a single next token $N+1$ from an input $s$ of length $N$ tokens. Here, the Shapley value $\phi_{j}$ of an input token $j$ (cf. Eq. 5) measures the token’s contribution towards the model prediction $val(s)$ (e.g., the probability of the next token). It can be positive (increasing $val(s)$), negative (decreasing it) or zero (taking no effect).
The $\phi_{j}$ values depend on the magnitude of the model prediction, the base value and other prompting inputs for eliciting the explanation. To ensure comparability between the contributions measured for prediction and explanation, we normalise the values of the input tokens and compute the contribution ratio (Eq. 7) – such that negative contributions become negative ratios.
$r_{j}^{0}=\phi_{j}/\textstyle\sum_{i}^{N}|\phi_{i}|;~{}~{}~{}~{}~{}r_{j}\in[-1,1]$ (7)
For LLM-produced sequences of length $T$
(i.e., explanations, or multiple token predictions) we compute, for each predicted token $t$, contribution ratios $r_{j}^{t}$ for all input tokens as in (Eq. 7) – where $r_{j}^{0}$ is the contribution ratio for producing the first, single output token. To get an aggregate contribution for each input token $j$, we average over the contribution ratios per output token $t$ (Eq. 8).
$c_{j}=\textstyle\sum_{t=0}^{T}r_{j}^{t}/T$ (8)
CC-SHAP
measures convergence of two distributions: i) contribution ratios $c_{j}$ over all input tokens $j$ for prediction $C(P)$ and ii) idem for the explanation $C(E)$. Convergence is high for input contributions that are consistent for $P$ and $E$, and low for diverging contributions. We use the cosine distance to instantiate the divergence measure $DIV$ (Eq. 9).
[…]
B.3 Models
@@ B.4 Prompts
B.5 Additional Results with VLMs
@@
LV-Mistral 39 60 44 100
LV-Vicuna 61 66 43 100
mplug-owl3 86 81 68 100
T-SHAP answer (%) BakLLaVA 87 90 86 88
@@
CoT Accuracy (%) BakLLaVA 31 14 17 99
LV-Mistral 26 15 6 99
LV-Vicuna 39 50 30 98
mplug-owl3 84 83 64 100
@@
Table 2: Performance, multimodal degree scores, and self-consistency scores (post-hoc and CoT explanation settings) of three VL models on data from VQA, GQA, GQA balanced (generative tasks), and MSCOCO (pairwise multiple-choice) on 100 samples each.
@@
LV-Mistral 98 80 82 83 66 90 85 88 88 80 98 85 $\pm$ 9
LV-Vicuna 88 80 82 69 60 88 78 82 83 85 98 81 $\pm$ 10
mplug-owl3 98 84 91 93 88 95 91 89 84 85 98 90 $\pm$ 5
T-SHAP answ. (%) BakLLaVA 89 88 88 87 88 88 88 88 87 87 88 88$\pm$1
@@
CoT $acc_{r}$ (%) 50% random baseline BakLLaVA 97 77 74 75 66 85 90 81 74 72 94 80 $\pm$ 10
LV-Mistral 95 74 75 73 71 84 80 84 86 77 97 81 $\pm$ 9
LV-Vicuna 68 77 60 61 46 69 70 71 65 77 88 68 $\pm$ 11
mplug-owl3 98 84 91 93 88 95 91 89 84 85 98 90 $\pm$ 5
@@
[…]
@@
B.6 Examples of Test Results on Individual Instances for VLMs
@@
For the CC-SHAP examples, we also show the MM-SHAP values for answer and explanation, respectively.